A 0.0290 M solution of an organic acid has an [H] of 0.00145 M. What is the pH of the solution, the percent ionization of the acid, and its Ka value? H= Number percent ionization Number K = Number
Answer #1

1)

[H+] = 1.45*10^-3 M

use:

pH = -log [H+]

= -log (1.45*10^-3)

= 2.84

Answer: 2.84

2)

let the weak acid be HA

HA -----> H+ + A-

2.9*10^-2 0 0

2.9*10^-2-x x x

x = [H+] = 1.45*10^-3 M

% ionization = x*100/c

= (1.45*10^-3)*100 / 0.0290

= 5.0 %

Answer: 5.0 %

3)

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Ka = 1.45*10^-3*1.45*10^-3/(0.029-1.45*10^-3)

Ka = 7.63*10^-5

Answer: 7.63*10^-5